OpenStudy (anonymous):

Find the mean, median, mode x|f 4|3 5|9 6|5 7|2 8|7 9|5 10|3 I tried to make the table lol:)

OpenStudy (anonymous):

I don't understand how to find the mean, median, mode with a table?

OpenStudy (anonymous):

mean is the average of all the frequency divide by 7 median u have to rearrange from ascending order and just find the middle number, if there's is two you have to add them together and divide by two. mode is the simply 5 because it has the most frequency and occurs the most

OpenStudy (anonymous):

I only look at the frequency numbers for median and mean?

OpenStudy (anonymous):

3 and 5 occur at the same time also for mode?

OpenStudy (anonymous):

well mode ur still looking at the highest occurring frequency still, which is number 5 clearly.

OpenStudy (anonymous):

the same amount I mean

OpenStudy (anonymous):

nvm im confused

OpenStudy (anonymous):

that has no impact on the mode at all.

OpenStudy (anonymous):

we are only looking at the highest frequency and we can see that is "9" at number 5.

jimthompson5910 (jim_thompson5910):

Mean = Sum(Value*Frequency)/(Total Frequency) Mean = (4*3+5*9+6*5+7*2+8*7+9*5+10*3)/(3+9+5+2+7+5+3) Mean = (12+45+30+14+56+45+30)/(3+9+5+2+7+5+3) Mean = (232)/(34) Mean = 6.8235

OpenStudy (anonymous):

ur right.

jimthompson5910 (jim_thompson5910):

No, there are 34 total values, not 7

jimthompson5910 (jim_thompson5910):

To find the median, start at the top and count up to 17 (this is half of 34) Then start at the bottom and count up to 17 You'll land at the same spot between x = 6 and x = 7 So the median is the average of these two: (6+7)/2 = 6.5